OpenStudy (anonymous):

guys please help i need to know how much 28 grams of stuff minus .8 in seprate baggies is how many baggies can you get if you ahve 28 g and only put .8 of a g in each bag

OpenStudy (akashdeepdeb):

I am assuming the question means this. You are given 28g of some stuff and you have to put 0.8g in bags? And you have to calculate the number of bags needed for that? ? Is that the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how many baggies of .8 can ya make

OpenStudy (akashdeepdeb):

 So 0.8g * (some number of bags) = 28g (some number of bags) = 28g/0.8g = 28/0.8 = 280/8 = 140/4 = 35. So you'll need 35 bags to equally distribute 28g into each one of them. Understood? :)